Calculate the area between the curves ? = ?2 and ? = 2? + 3 Hint: Use algebra or Desmos or a quick sketch of graph to determine the limits of integration.
Answer #1

x² = 29+3 2 a_27-3=0 Upper limit it) Sol:- y=x²; y=2a+3 as-1s 0-3 1st Gump ਤੇ ਨਾ Hy=x² as well 5 Y=2x+3 125 PS, BET ya -1 3 وplease comment if you have any doubts will clarify & pls upvote if it helps